OpenStudy (anonymous):

how to express 5 in an equivalent form?

OpenStudy (kainui):

Not sure what you're looking for, something like 10/2?

OpenStudy (anonymous):

im not really sure how to do this

OpenStudy (mathstudent55):

Use an equivalent fraction: 5 = 5/1 = 10/2 = 15/3 = ...

OpenStudy (anonymous):

thanks